OpenStudy (anonymous):

Given vectors a and b find: (a)2a (b)2a+3b (c)b-3a 1) a=2i, b=i+j

OpenStudy (anonymous):

part (a) 2(2i) = 4i, part (b) 4i +3i+3j = 7i+3j, part (c) i+j-3i=-2i+j

OpenStudy (anonymous):

How did you get part a and b?

OpenStudy (anonymous):

because a = 2i, and you are asked to multiply 2 by 2i so you get 4i

OpenStudy (anonymous):

for part b you already know from part a that 2a is 4i, and you were asked to to multuply b which is i + j by 3, so you get 3i + 3j, then you add 2a + 3b you get the above result...

OpenStudy (anonymous):

2(2i) + 3(i+j)= 4i +3i+3, you add the i components, so you get 7i + 3j

OpenStudy (anonymous):

Oh ok thank you

OpenStudy (anonymous):

you are welcome!